# InkSight: Offline-to-Online Handwriting Conversion by Learning to Read and Write

### v2

@@ Abstract
Digital note-taking is gaining popularity, offering a durable, editable, and easily indexable way of storing notes in the vectorized form, known as digital ink. However, a substantial gap remains between this way of note-taking and traditional pen-and-paper note-taking, a practice still favored by a vast majority. Our work, InkSight, aims to bridge the gap by empowering physical note-takers to effortlessly convert their work (offline handwriting) to digital ink (online handwriting), a process we refer to as derendering. Prior research on the topic has focused on the geometric properties of images, resulting in limited generalization beyond their training domains. Our approach combines reading and writing priors, allowing training a model in the absence of large amounts of paired samples, which are difficult to obtain. To our knowledge, this is the first work that effectively derenders handwritten text in arbitrary photos with diverse visual characteristics and backgrounds. Furthermore, it generalizes beyond its training domain into simple sketches. Our human evaluation reveals that 87% of the samples produced by our model on the challenging HierText dataset are considered as a valid tracing of the input image and 67% look like a pen trajectory traced by a human. Interactive visualizations of 100 word-level model outputs for each of the three public datasets are available in our Hugging Face space. Model release is in progress.
Machine Learning, ICML
Figure 1: Results of InkSight. Left: Image of handwriting (offline handwriting), Right: output digital inks (online handwriting). In every word, character colors transition from red to purple, following the rainbow sequence; within each stroke, the shade progresses from darker to lighter. More results in Appendix A.
@@ 1 Introduction
Handwritten notes have been a cornerstone of information storage for centuries. Today, with the rise of stylus and digital pen technologies, digital inking presents a compelling alternative. This modern approach offers several advantages: enhanced durability, seamless organization and integration with other digital content (images, text, links) or digital assistance. Despite these benefits, many people still prefer traditional handwritten notes over the digital format.
Our work aims to make physical notes, particularly handwritten text, available in the form of digital ink, capturing the stroke-level trajectory details of handwriting. This allows paper note-takers to enjoy the benefits of digital medium without the need to use a stylus. This area has gained significant interest in both academia (Nguyen et al., 2021; Chen et al., 2022b; Mohamed Moussa et al., 2023) and industry (Not, 2023), with software solutions that digitize handwriting and hardware solutions that require smart pens and/or special paper (Liv, 2024; Roc, 2024; Neo, 2024). Our approach requires only a picture of the handwritten note, without the need for specialized equipment.
Our approach differs from the methods that rely on geometric priors, where gradients, contours, and shapes in an image are utilized to extract writing strokes. Instead, we harness the power of learned reading and writing priors, where:
  • Learned reading prior enhances the model’s capability in precisely locating and extracting textual elements from the images. This is achieved either through the model’s textual understanding ability or aided with external textual input, e.g. from an OCR engine.
  • The integration of writing prior ensures that the resulting vector representation, the digital ink, closely aligns with the typical human approach of writing in terms of physical dynamics and the order of strokes.
To the best of our knowledge, our work is the first to incorporate these priors, resulting in a robust model that is capable of derendering handwriting across diverse scenarios and appearances, including challenging lighting conditions, noticeable occlusions, etc.
Our model adopts a simple architecture combining the widely popular and readily available ViT (Dosovitskiy et al., 2021) encoder and an mT5 (Xue et al., 2021) encoder-decoder, fostering reproducibility, reusability, and ease of adoption.
To summarize, the major contributions of our work are:
@@
    We propose the first system to perform derendering, transforming arbitrary photos of handwritten text into digital ink.
@@
    We propose a training and inference setup that works without expensive data collections and scales to arbitrarily large input images.
@@
    We show that the inks produced by our system are both semantically and geometrically similar to the input images, and are similar to real digital ink data, as measured by both automatic and human evaluations.
@@
    We show that, due to the learned reading and writing priors, our approach is robust well beyond its training data and works on various types of handwriting, simple sketches, and full pages of notes.
@@ 2 Related Work
Pen trajectory recovery
has been a task of interest to many researchers due to its utility for tasks that can benefit from stroke order / temporal information such as online handwriting recognition (Lallican et al., 2004; Viard-Gaudin et al., 2005; Zhang et al., 2015; Chan, 2020). Another point of interest is the ability to efficiently store, index, and edit handwritten notes within existing online note taking systems.
Classical approaches commonly consist of domain-specific preprocessing (e.g. noise removal, image binarization, skeletonization), local (sub-)stroke level processing (e.g. identification of junction points) and global aggregation (usually as a graph-based optimization problem) (Jager, 1996; Kato & Yasuhara, 2000; Qiao et al., 2006; Chan, 2020; Doermann et al., 2002). These approaches often rely on the quality of the preprocessing and hand-designed heuristics and do not generalize well to other scripts and domains.
More recent methods use convolutional (Nguyen et al., 2021; Archibald et al., 2021) and/or recurrent (Bhunia et al., 2018; Chen et al., 2022b) neural networks to translate an image into a sequence of coordinates. Mohamed Moussa et al. (2023) use two Transformer models to first encode the substrokes of the input and then reorder them, given the encodings. Those methods produce promising results, but focus on the simplified setups of rendered online handwriting and/or single characters.
Figure 2: Left: Illustration of model inference. Inputs include an image and a prompt specifying the task. Outputs may consist of ink alone or a combination of ink and text, as indicated by matching colors in both the input prompt and the output. Ink tokens are detokenized into digital ink using our proposed ink tokenizer explained in Section 3.3. Right: Diagram of the Full Page System.
Line drawing vectorization
shares a lot of features with geometric approaches of pen trajectory recovery, but does not include stroke order reconstruction.
Recent techniques for this problem involve solving an optimization task of fitting a set of geometric primitives (e.g. Bezier curves) to match the geometric and/or semantic content of the input image (e.g. Vinker et al. 2022), sometimes relying on differentiable rendering to maintain the inputs and outputs in the image space Mo et al. (2021).
@@
These approaches focus on converting raster images (perfectly aligned, clean, clear backgrounds) to vector ones. Thus they avoid dealing with the realistic photo artifacts related to lighting, noisy backgrounds or photo-taking angles. These artifacts are however unavoidable in a real-world setting.
Dataset availability
@@
is limiting the research in trajectory recovery as there are few datasets of images and their corresponding digital inks. These include IRONOFF (Viard-Gaudin et al., 1999), CASIA (Liu et al., 2011), IBM-UB (Shivram et al., 2013) for handwritten text (images there exhibit limited variability being always black-on-white line writing, usually on the same input device) and Sketchy (Sangkloy et al., 2016) for crowd-sourced sketches of photos from 125 object categories.
Combining language models with new modalities
is related to this work as we convert the pen trajectory to a sequence of coordinates on a fixed size canvas and represent the coordinates with ”ink tokens” (more details in Section 3.3). Hence, we consider our approach as extending a Vision-Language Model (VLM) with a new modality.
@@
Recent research in this area has been pointed towards training adapters for merging the pretrained unimodal models (e.g. Alayrac et al. (2022)) or, fully or partially fine-tuning a system with multi-modal tasks (e.g. Chen et al. (2022a)). Works most similar to ours are AudioPaLM (Rubenstein et al., 2023), extending LLM with audio tokens, Painter (Pourreza et al., 2023) using coordinate-based representation of strokes to perform generation and understanding of sketches, and PixelLLM (Xu et al., 2023) using the same representation to localize words in the image caption.
@@ 3 Method
While the fundamental concept of InkSight appears straightforward – training a model that generates digital ink representations from input images – the practical implementation for arbitrary input images presents two significant challenges: (1) Limited Supervised Data: acquiring paired data with corresponding images and ground truth digital ink for supervised training can be expensive and time-consuming and no datasets with sufficient diversity exist for this task. (2) Scalability to Large Images: the model must effectively handle potentially arbitrary large input images with varying resolutions and amount of content.
To address the first problem without expensive data collection, we propose a multi-task training setup which combines recognition and derendering tasks. We show that this multi-task training setup enables the model to generalize on derendering tasks with various styles of images as input, and injects the model with both semantic understanding and writing priors of handwritten text. We discuss our data and multi-task training setup in Section 3.2, and ablate the design choices in Section 4.6. The exact model structure, which consists of standard publicly available off-the-shelf components, is discussed in Section 3.1.
The second problem can be addressed by training a model with very high-resolution input images and very long output sequences, but this is computationally prohibitive. Instead, we break down the derendering of a page of notes into three steps: running OCR to extract word-level bounding boxes, derendering each of the words separately and pasting the derendered words back, as shown in Figure 2. We discuss the representation of digital ink, the normalization of the data, and the tokenization scheme we use in Section 3.3. More full page results can be found in Appendix L.
@@
3.1 Vision-Language Model for Digital Ink
In this study, we employ a model architecture inspired by PaLI (Chen et al., 2022a, 2023b, 2023a) which integrates a Vision Transformer (ViT) encoder (Dosovitskiy et al., 2021) with an mT5 encoder-decoder Transformer model (Xue et al., 2021). We provide the task-specific instructions as text, as described in Section 3.2. We use the same set of tokens for both input and output, which contains the individual character tokens from the original mT5 vocabulary and specific tokens reserved for the ink representation (further details in Sec. 3.3).
To initialize the model, we use the pre-trained weights of the ViT encoder. The mT5-based encoder-decoder weights are initialized randomly, which is motivated by our customized token dictionary differing from the one used in mT5’s original training. In our training, we freeze the weights of the ViT encoder – we justify this choice empirically in Section 4.6.
@@ 3.2 Training Task Mixture
Tasks Description
Vanilla Derender
The model receives a synthetic ink image with a conditioning task prompt and outputs ink tokens corresponding to the image content.
Derender with Text
The model receives a synthetic ink image, the target textual content together with a task conditioning prompt, and outputs ink tokens corresponding to the provided text within the image.
Recognition (Syn/Real)
The model receives an ink image (either synthetic or real) and a task conditioning prompt and outputs the recognized textual content. This task uses both synthetic ink images and real images from OCR datasets.
Recognize and Derender
This task combines recognition and derendering. The model receives a synthetic ink image and a prompt and outputs both the recognized text and the corresponding ink tokens.
We show that this setup helps the model to (1) generalize to the derendering of real photos; (2) learn useful priors which help dealing with occlusions and generating realistic ink; (3) allow using different inference setups which enables, e.g. high-quality text derendering with OCR result (Derender with Text) or derendering without textual input for non-textual elements (Vanilla Derender).
Training tasks are shuffled and assigned equal appearance probability (20%). By default, during inference, we use Derender with Text (this choice is ablated in Section 4.6).
@@
To narrow the domain gap between the synthetic images of rendered inks and the real photos, we augment the data in tasks that take rendered ink as input. This is done by randomizing the ink angle, color, stroke width, and by adding Gaussian noise and cluttered backgrounds. Examples are provided in Figure 3, more details are given in Appendix E, and Section 4.6 highlights the importance of having the data augmentation.
@@ 3.3 Data Representation
Digital ink is usually represented as a sequence of strokes $I=\{s_{1},s_{2},\cdots,s_{n}\}$, where each stroke $s_{i}$ consists of a sequence of $m_{i}$ (representing $i$-th stroke length) coordinate-time triplets, denoted as $s_{i}=\left\{(x_{i},y_{i},t_{i})\right\}_{i=1}^{m_{i}}$. To enable text decoders like mT5 to generate ink strokes, we first normalize the ink and then tokenize it into a set of discrete tokens from a pre-defined dictionary.
Figure 4: Illustration of the ink tokenization for a single-stroke ink. The dark red ink depicts the normalized ink stroke, with numbered circles marking sampled points after time resampling. The color gradient of the sampled points indicates the point order. Each point is represented with two tokens encoding its $x$ and $y$ coordinates. The token sequence for this ink begins with b, signifying the start of the stroke, followed by the tokens for coordinates of sampled points.
Normalizing the ink includes (1) resampling it at a fixed frequency (20 ms) to account for potential differences between input devices; (2) applying Ramer-Douglas-Peucker resampling (Visvalingam & Whyatt, 1990) to reduce the output sequence length while preserving the overall shape; (3) shifting and scaling the ink to place it at the center of fixed size canvas, such that each point in each stroke is within the range of $[0,N]$.
@@
For the derendering tasks of the training mixture, we render digital inks in the center of an $M\times M$ image with stroke width, color, and background color selected via random augmentation (see Appendix E). For the recognition task training and inference with real images, we apply a similar transformation: the input image is scaled and then centered and padded in black to produce an $M\times M$ image, although we use $M=224$ in practice, by design there are no constraints to have $N=M$.
@@
Our set of tokens contains all individual symbols from the multilingual mT5 tokenizer (Xue et al., 2021) (appx. 20k), with additional $32N+$ 3 tokens reserved for the representation of the digital ink, as described above. Removing all multi-symbol tokens allows to reduce the size of the input text embedding and final softmax of the decoder model by $\sim$ 80%, while maintaining the ability to input and output arbitrary text and not affecting the quality of derendering or recognition.
4 Results

### v4

# InkSight: Offline-to-Online Handwriting Conversion by Teaching Vision-Language Models to Read and Write

@@ Abstract
Digital note-taking is gaining popularity, offering a durable, editable, and easily indexable way of storing notes in a vectorized form, known as digital ink. However, a substantial gap remains between this way of note-taking and traditional pen-and-paper note-taking, a practice that is still favored by a vast majority. Our work InkSight [^1] GitHub: https://github.com/google-research/inksight Hugging Face: [link] Project page: [link] , aims to bridge the gap by empowering physical note-takers to effortlessly convert their work (offline handwriting) to digital ink (online handwriting), a process we refer to as derendering. Prior research on the topic has focused on the geometric properties of images, resulting in limited generalization beyond their training domains. Our approach leverages vision-language models to combine reading and writing priors, allowing training a model in the absence of large amounts of paired samples, which are difficult to obtain. To our knowledge, this is the first work that effectively derenders handwritten text in arbitrary photos with diverse visual characteristics and backgrounds. Furthermore, it generalizes beyond its training domain into simple sketches. Our human evaluation reveals that 87% of the samples produced by our model on the challenging HierText dataset are considered as a valid tracing of the input image and 67% look like a pen trajectory traced by a human.
Figure 1: Results of InkSight [Animated visualization]. Left: A photo of handwritten text (offline handwriting), Right: Output digital ink (online handwriting). In every word, character colors transition from red to purple, following the rainbow sequence, ROYGBIV, which reflects the stroke order; within each stroke, the shade progresses from darker to lighter. More visualizations in Appendix B.
@@ 1 Introduction
Handwritten notes have been a cornerstone of information storage for centuries. Today, with the rise of stylus and digital pen technologies, digital inking is a compelling alternative. This modern approach to storing information offers several advantages: Digital ink offers enhanced durability, structured organization, and fine-grained editability. Unlike pen-and-paper notes, it supports post hoc modifications users can move, resize, or restyle individual strokes or sections. This enables dynamic reuse, seamless integration with digital workflows, and compatibility with assistive systems. Despite these apparent benefits, many people still prefer traditional handwritten notes over the digital format.
Traditional OCR extracts textual content from handwriting but discards the rich structural and temporal information embedded in the handwritten form. In contrast, our approach captures handwriting as digital ink, preserving stroke-level trajectories. This enables users to retain the familiarity of pen-and-paper note-taking while gaining the flexibility of digital representations without requiring a stylus.
The topic of converting offline handwriting to online handwriting has gained significant interest in both academia (Nguyen et al., 2021; Chen et al., 2022b; Mohamed Moussa et al., 2023) and industry (Not, 2023), with the introduction of software solutions that digitize handwriting and hardware solutions that require smart pens and/or special gadgets  (Nuw, 2025; Liv, 2024; Roc, 2024; Neo, 2024). Our approach requires only an image of the handwritten note, without any specialized equipment. It differs from the methods that rely on geometric priors, where gradients, contours, and shapes in an image are utilized to extract writing strokes. Instead, we harness the power of learned reading and writing priors, where:
  • The learned reading prior enhances the model’s capability to precisely locate and extract textual elements from images. This is achieved either through the model’s textual understanding ability or aided with external textual input, e.g. from an OCR engine.
  • The integration of the writing prior ensures that the resulting vector representation, the digital ink, closely aligns with the typical human approach to writing in terms of physical dynamics and the order of strokes.
To the best of our knowledge, our work is the first to incorporate these priors, resulting in a robust model that is capable of derendering handwriting across diverse scenarios and appearances, including challenging lighting conditions, noticeable occlusions, etc. Our model adopts a simple architecture combining the widely popular and readily available ViT (Dosovitskiy et al., 2021) encoder and an mT5 (Xue et al., 2021) encoder-decoder, fostering reproducibility, reusability, and ease of adoption. To summarize, the major contributions of our work are:
@@
    We propose the first system to perform derendering: transforming arbitrary photos of handwritten text into digital ink. It relies only on the standard architecture components and the training and inference setup that works without expensive data collections and scales to arbitrarily large input images.
@@
    We show that the inks produced by our system are both semantically and geometrically similar to the input images. We demonstrate that they are similar to real digital ink data, as measured by both automatic and human evaluations.
@@
    We show that, with the learned reading and writing priors, our approach is robust and works on various types of handwriting, simple sketches, and full pages of notes.
@@
    We release the public model Small-p, trained entirely on publicly available datasets, along with a subset of model-generated inks in universal inkML format and expert-traced online handwriting data to support future research on the topic.
@@ 2 Related Work
Pen trajectory recovery has been a task of interest to many researchers due to its utility for tasks that can benefit from stroke order / temporal information such as online handwriting recognition (Lallican et al., 2004; Viard-Gaudin et al., 2005; Zhang et al., 2015; Chan, 2020). Another point of interest is the ability to efficiently store, index, and edit handwritten notes within existing online note-taking systems. Classical approaches commonly consist of domain-specific preprocessing (e.g. noise removal, image binarization, skeletonization), local (sub-)stroke level processing (e.g. identification of junction points) and global aggregation (commonly as a graph-based optimization problem) (Jager, 1996; Kato & Yasuhara, 2000; Qiao et al., 2006; Chan, 2020; Doermann et al., 2002). These approaches often rely on the quality of the preprocessing and hand-designed heuristics and do not generalize well to other scripts and domains. More recent methods use convolutional (Nguyen et al., 2021; Archibald et al., 2021) and/or recurrent (Bhunia et al., 2018; Chen et al., 2022b) neural networks to translate an image into a sequence of coordinates. Set, sort! (Mohamed Moussa et al., 2023) use two Transformer models to first encode the substrokes of the input and then reorder them, given the encodings. Those methods produce promising results, but focus on the simplified setups of rendered online handwriting and/or single characters.
@@
Line drawing vectorization shares many features with geometric approaches to pen trajectory recovery, but does not include stroke order reconstruction. Recent techniques for this problem involve solving an optimization task of fitting a set of geometric primitives (e.g. Bezier curves) to match the geometric and/or semantic content of the input image (Vinker et al., 2022), sometimes relying on differentiable rendering to maintain the inputs and outputs in the image space (Mo et al., 2021). These approaches focus on converting raster images (perfectly aligned, clean, clear backgrounds) to vector ones. Thus they avoid dealing with the realistic photo artifacts related to lighting, noisy backgrounds or photo-taking angles. These artifacts are however unavoidable in a real-world setting.
@@
Dataset availability is limiting the research in trajectory recovery as there are few datasets of images and their corresponding digital inks. These include IRONOFF (Viard-Gaudin et al., 1999), CASIA (Liu et al., 2011), IBM-UB (Shivram et al., 2013) for handwritten text (images there exhibit limited variability being always black-on-white line writing, usually on the same input device) and Sketchy (Sangkloy et al., 2016) for crowd-sourced sketches of photos from 125 object categories.
@@
Combining language models with new modalities is related to this work since we convert the pen trajectory to a sequence of coordinates on a fixed size canvas and represent the coordinates with ink tokens (more details in Section 3.3). Hence, we consider our approach as extending a Visual Language Model (VLM) with a new modality. Recent research in this area has been pointed towards training adapters for merging the pretrained unimodal models (e.g. Alayrac et al. (2022)) or, fully or partially fine-tuning a system with multi-modal tasks (e.g. Chen et al. (2022a)). Works most similar to ours are AudioPaLM (Rubenstein et al., 2023), extending LLM with audio tokens, Painter (Pourreza et al., 2023) using coordinate-based representation of strokes to perform generation and understanding of sketches, and PixelLLM (Xu et al., 2024) using the same representation to localize words in the image caption.
@@ 3 Method
While the fundamental concept of derendering appears straightforward – training a model that generates digital ink representations from input images – the practical implementation for arbitrary input images presents two significant challenges:
  1. 1.
    Limited Supervised Data: Acquiring paired data with corresponding images and ground truth digital ink for supervised training can be expensive and time-consuming and no datasets with sufficient diversity exist for this task yet.
  2. 2.
    Scalability to Large Images: The model must effectively handle potentially arbitrary large input images with varying resolutions and amounts of content.
Figure 2: Left [Animated visualization]: Illustration of model inference. Inputs include an image and a prompt specifying the task. Outputs may consist of ink alone or a combination of ink and text, as indicated by matching colors in both the input prompt and the output. Ink tokens are detokenized into digital ink using our proposed ink tokenizer explained in Section 3.3. Right [Animated visualization]: Diagram of the Full Page System.
To address the first problem without expensive data collection, we propose a multi-task training setup that combines recognition and derendering tasks. We show that this setup enables the model to generalize on derendering tasks with various styles of images as input, and injects the model with both semantic understanding and writing priors of handwritten text. We discuss our data and multi-task training setup in Section 3.2, and ablate the design choices in Section 4.6. The exact model structure, which consists of standard publicly available off-the-shelf components, is discussed in Section 3.1. In Section 3.3, we cover the representation of digital ink, the normalization of the data, and the tokenization scheme we employ.
The second problem can be addressed by training a model with very high-resolution input images and very long output sequences, but this is computationally prohibitive. Instead, we break down the derendering of a page of notes into three steps: running OCR to extract word-level bounding boxes, derendering each of the words separately and pasting the derendered words back, as shown in Figure 2. An example of full page inference is shown in Figure 1, more full page results are presented in Appendix N.
3.1 Vision-Language Model for Digital Ink
Our model architecture, inspired by that of PaLI (Chen et al., 2022a; 2023b; 2023a), integrates a Vision Transformer (ViT) encoder (Dosovitskiy et al., 2021) with an mT5 encoder-decoder Transformer model (Xue et al., 2021). We provide the task-specific instructions as text, as described in Section 3.2. We use the same set of tokens for both input and output, which contains the individual character tokens from the original mT5 vocabulary and specific tokens reserved for the ink representation (details in Sec. 3.3). To initialize the model, we use the pre-trained weights of the ViT encoder. The mT5-based encoder-decoder weights are initialized randomly, which is motivated by our customized token dictionary differing from the one used in mT5’s original training. In our training, we freeze the weights of the ViT encoder – we justify this choice empirically in Section 4.6.
@@ 3.2 Training Task Mixture
Task Description
Vanilla Derender The model receives a synthetic ink image with a conditioning task prompt and outputs ink tokens corresponding to the image content.
Derender with Text The model receives a synthetic ink image and a task conditioning prompt including the target textual content, and outputs ink tokens corresponding to the provided text within the image.
Recognition (Syn/Real) The model receives a handwriting image (either synthetic or real) and a task conditioning prompt and outputs the recognized textual content. This task uses both synthetic ink images and real images from OCR datasets.
Recognize and Derender This task combines recognition and derendering. The model receives a synthetic ink image and a prompt and outputs both the recognized text and the corresponding ink tokens.
We show that this setup helps the model to: (1) generalize to the derendering of real photos; (2) learn useful priors that help deal with occlusions and generate realistic ink; (3) allow using different inference setups that enable, e.g. high-quality text derendering with OCR result (Derender with Text) or derendering without textual input for non-textual elements (Vanilla Derender). Training tasks are shuffled and assigned equal appearance probability (20%). However, future work could explore adjusting this ratio dynamically during training and towards the end of training. For best performance during inference, we use Derender with Text (this choice is ablated in Section 4.6 where we show that our model performs well without reliance on OCR as well).
@@
To narrow the domain gap between the synthetic images of rendered inks and the real photos, we augment the data in tasks that take rendered ink as input. This is done by randomizing the ink angle, color, stroke width, and by adding Gaussian noise and cluttered backgrounds. Examples are provided in Figure 3, more details are given in Appendix F, and Section 4.6 highlights the importance of having the data augmentation.
@@ 3.3 Data representation
Digital ink is usually represented as a sequence of strokes $I=\{s_{1},s_{2},\cdots,s_{n}\}$, where each stroke $s_{i}$ consists of a sequence of $m_{i}$ (length of the $i$-th stroke) coordinate-time triplets, represented as $s_{i}=\left\{(x_{i},y_{i},t_{i})\right\}_{i=1}^{m_{i}}$. To enable sequence models like mT5 to generate ink strokes, we first normalize the ink and then tokenize it into a set of discrete tokens from a pre-defined dictionary. Normalizing the ink includes:
  1. 1.
    Resampling: Resample the ink at a fixed frequency (20 ms) to account for potential variations between input devices.
  2. 2.
    Sequence Reduction: Apply Ramer-Douglas-Peucker resampling (Visvalingam & Whyatt, 1990) to reduce the sequence length while preserving the overall shape of the strokes.
  3. 3.
    Normalization: Shift and scale the ink strokes to position them at the center of a fixed-size canvas. Each point in a stroke is scaled to fit within the range $[0,N]$, where we use $N=224$ in our implementation.
Figure 4: Ink tokenization for a single-stroke digital ink [Animated visualization]. The dark red ink depicts the normalized ink stroke, with numbered circles marking sampled points after time resampling. The color gradient of the sampled points indicates the point order. Each point is represented with two tokens encoding its $x$ and $y$ coordinates. The token sequence for this ink begins with b, signifying the beginning of the stroke, followed by the tokens for coordinates of sampled points.
@@
For the derendering tasks of the training mixture, we render digital inks in the center of an $M\times M$ image with stroke width, color, and background color selected via random augmentation (see Appendix F). For the recognition task training and inference with real images, we apply a similar transformation: the input image is scaled, centered, and padded in black to produce an $M\times M$ image. Although we use $M=224$, by design there are no constraints to have $N=M$.
@@
Our set of tokens contains all individual symbols from the multilingual mT5 tokenizer (Xue et al., 2021) (approx. 20k), with additional $32N+$ 3 tokens reserved for the representation of the digital ink, as described above. Removing all multi-symbol tokens allows to reduce the size of the input text embedding and final softmax of the decoder model by $\sim$ 80%, while maintaining the ability to input and output arbitrary text and not affecting the quality of derendering or recognition.
4 Experiments